Calculate the molar mass of the following substances.
       (a) Ethyne, C2H2
       

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

Ethyne, C2H2= 2 × 12 + 2 × 1 = 24 + 2 = 26 g
Thanks